OpenStudy (anonymous):

A road has a 10% grade. a.) What is the angle of elevation of the road? b.) If the road is 2 km. long, how much does it rise?

OpenStudy (lastdaywork):

Do you know what does "A road has a 10% grade." imply??

OpenStudy (anonymous):

no, it just says to find the angle of elevation probably using sin cosin or tangent

OpenStudy (lastdaywork):

Read - http://en.wikipedia.org/wiki/Grade_%28slope%29 Now can you tell me the angle of elevation ??

OpenStudy (lastdaywork):

Actually, just see the equations - http://en.wikipedia.org/wiki/Grade_%28slope%29#Equations

OpenStudy (lastdaywork):

Okay..the relation between the grade and angle of elevation is as follows: Grade (in %) = tan(x) * 100 Now can you tell me the value of x ??

OpenStudy (anonymous):

no, the way i was taught was using triangles and pictures but this problem doesn't have a picture and my notes have nothing about anything with a grade

OpenStudy (campbell_st):

here is a simple concept... 10% grade... so it rises 1 for every 10 that the road runs  since 1/10 = 10% so tan(x) = 1/10 easy to find x now

OpenStudy (anonymous):

5.7?

OpenStudy (campbell_st):

ok... so the angle is correct... so for part (b) the question is now  so you now need cos to find a sin and b hope it helps

OpenStudy (campbell_st):

you may need to change 2km to 2000 metres

OpenStudy (anonymous):

a=1990 b=199

OpenStudy (anonymous):

troroorororrorolololo people hello fron ua

OpenStudy (campbell_st):

well not quite for a you need cos cos(5.7) = a/2000 so a = 2000*cos(5.7) and for b its sin sin(5.7) = b/2000 b = 2000*sin(5.7) hope it helps

OpenStudy (anonymous):

troroorororrorolololo people hello fron ua

OpenStudy (anonymous):

its trololololololololololololololololololololo not trorororororororororororororo pavlo fail picasso

OpenStudy (anonymous):

pikagirl

OpenStudy (anonymous):

its pikachu

OpenStudy (anonymous):

nonon it is pikagirl

OpenStudy (anonymous):

nononononono you no understand, its pikachu not pikagirl you pavlo fail picasso

OpenStudy (anonymous):

i have no idea thanks anyways @campbell_st

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

you're welcome

OpenStudy (campbell_st):

@ileah I just checked your answers and they are correct... sorry for the misinformation

OpenStudy (anonymous):

you mader traker i will kill you

OpenStudy (anonymous):

thanks (: @campbell_st but im still not sure what the answer would be. i need to know how much it will rise

OpenStudy (campbell_st):

well you posted 1990 for the run and 199 for the rise they are correct

OpenStudy (anonymous):

okay(: @campbell_st would you mind helping me with another problem?

OpenStudy (campbell_st):

Sorry I have to go... there are plenty of people who can help... just page some of the high number people.

OpenStudy (anonymous):

thanks